OpenStudy (anonymous):

Solve and Check 4(2x-3)=3(2x+6)

OpenStudy (anonymous):

4(2x-3)=3(2x+6) 8x-12=6x+18 2x=30 x=15

OpenStudy (anonymous):

thank u so much